OpenStudy (anonymous):

Suppose a machine has five independent components each of which has probability 0.01 of failing , (a) find the probability that none or all components will fail, (b) find the probability that at least two components will fail?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

all fail: \((0.01)^5\)

OpenStudy (anonymous):

none fail \((.99)^5\)

OpenStudy (anonymous):

as the question is finding all fail or none fail, i should add (0.01)5 and (0.99)5

OpenStudy (anonymous):

oh i thought that was two different numbers yes, since they are obviously mutually exclusive, add them up

OpenStudy (anonymous):

oh, thank you, could you help me in finding the next part in the question i.e., finding probability that at least two components will fail

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

"at least two" translates as 2 or 3 or 4 or 5 which is a lot to compute much easier to think of "at least 2" as "not 0, not 1" compute those, subtract from 1

OpenStudy (anonymous):

you already got 0, what was \(.99^5\)

OpenStudy (anonymous):

do you know how to get 1?

OpenStudy (anonymous):

could you tell me how to get for 1?

OpenStudy (anonymous):

sure one fails \(.01\) 4 do not fail \(.99^4\) and there are 5 ways for this to happen, first one fails others don't second one fails, others don't, etc

OpenStudy (anonymous):

so \[5\times (0.1)\times (.99)^4\]

OpenStudy (anonymous):

it is the binomial probability with \(p=.99,1-p=.01, n=5\) general form is \[P(x=k)=\binom{n}{k}p^k(1-p)^{n-k}\]

OpenStudy (anonymous):

oh Thank You very much:>

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

hey hi, could you solve this question for me A circuit consists of 8 resistors, 2 inductors and 3 capacitors. 3 of the circuit elements are damaged. Find the following probabilities (a) p(only 1 capacitor is damaged) (b) p(at least one inductor is damaged) (c) p(no resistor are damaged).